# HARMONIOUS – Human-like reactive motion control and multimodal perception for humanoid robots

### v1

@@ III-C Real robot – bimanual task
III-D Real robot – Interactive game demonstration
To demonstrate the operation of HARMONIOUS at scale and in a genuine unstructured human-robot interaction scenario, we taught the robot to play a children’s board game called Bubbles [^1] A description of the game is available at https://www.piatnik.com/en/games/board-games/children-games/bubbles. For our purposes, we simplified the game, with only six cards, three dice, only the human rollings the dice, and without grasping the cards (only pointing at the selected one). The robot recognizes the cards but is told the number of the dice after the roll (as these are too small to be recognized).. Based on the numbers that appear on the dice and their colors, the players—the robot and the human in our case—compete to pick the card on the table where the size of the colored bubbles corresponds to the colors of the dice, ordered by the numbers – see Fig. 1 and a video at https://youtu.be/gw8JB-1R3bs. This game was chosen because there is a shared physical space with unstructured interaction. The robot processes the visual inputs and recognizes individual cards on the table. After the roll of the dice, the player that first places his hand over the right card wins the round. After choosing the appropriate card, the robot selects which hand to use and initiates the reaching movement. At all times and in real time, visual, proximity and tactile streams are simultaneously processed, spawning obstacles if appropriate and dynamically adding constraints to the robot controller in order to warrant the safety of the human player at all times. The game itself is rather challenging for the human player and hence in order to compete with the robot, he often cheats by pushing the robot physically away or by blocking certain volumes of the workspace by placing his hands over it. This was a deliberate choice to put HARMONIOUS at test.
The controller runs in the dual-arm mode during the game. Depending on the position of the card in every round, the left or right arm is chosen and set as “primary’ (and hence its position reaching target is set as a constraint). Additional constraints were added to the optimization problem by adding static obstacle points (see Sec. II-I) to prevent the robot from colliding with the table. We prepared a high-level Python script to play the game and used the gaze controller [53] to always look at the table. The program loop starts with the processing of rolled dice (numbers sent to the program in the command line) and the current RGB-D image, followed by the detection of cards on the table, the decision of the correct card and computing its 3D position. Then one arm approaches the card position; while the other arm goes to a specific position outside the field of view of the RGB-D camera. After the reach approach (successful or not), the approaching arm returns to a specific position to clear the view for the next round.
@@ III-D Real robot – Interactive game demonstration
IV Conclusion, Discussion and Future work
@@
We have developed a real-time reactive motion control system for upper body control of a dual-arm robot with a common torso. The proposed solution (HARMONIOUS) is designed to enable safe and effective interaction with humans in close proximity. In the absence of obstacles, we systematically tested our controller’s reaching ability on a grid of position and orientation targets in the workspace and compared better or at least equally good as alternative controllers, including the one designed specifically for the iCub [4]. We found that assigning different weights to position and orientation tasks is advantageous (unlike in [8]). Preferring arm over torso movements by adjusting weights in the minimization leads to more natural-looking movements. Kinematic singularities are handled through velocity damping and natural kinematic postures are rewarded in the optimization problem formulation. Smooth human-like motions with a bell-shaped Cartesian velocity profile are generated, using biologically inspired local trajectory sampling. Overall, several components act synergestically to produce human-like naturally looking movements.
The full potential of HARMONIOUS lies in the ability to perceive objects in the environment, relate them to the complete surface of the robot body, and, when appropriate, transform them into constraints for the motion controller. We used three sensory modalities: two that sense at a distance, “pre-collision”, vision and proximity, and tactile sensors on large areas of the robot body (2000 individual sensors) providing contact or post-collision information. Information about obstacles from all three modalities—sometimes overlapping, at other times complementary—was processed and transformed into a unified representation dynamically generating constraints for the whole-body motion controller. We have extensively experimentally evaluated HARMONIOUS, compared it to NEO [8] and React [17, 30], demonstrating superior performance. Uniquely to our solution, we commanded both arms and the torso of the iCub humanoid robot (in total 17 DoFs) and demonstrated both the setup where every arm has a separate task and one, set as “primary arm”, commands the torso joints, and a bimanual task where the hands hold an object between them. The performance of HARMONIOUS at scale was finally demonstrated in unstructured physical human-robot interaction while playing a game. Here, additional task constraints were added to the problem formulation.
The solution we presented and its experimental demonstrations are unique in showing the online dynamic incorporation of obstacles perceived through three different sensory modalities and their on the run embedding as whole-body motion constraints into a motion controller with human-like characteristics. At the same time, our solution is by no means restricted to the iCub humanoid robot or to the particular sensory setup. The problem formulation is highly modular and both the minimization criteria (e.g. rewarding natural postures) or constraints to the quadratic program (e.g. obstacles) can be easily removed or different ones added.
@@
Our immediate future work involves adding active gaze control, such that the robot can better perceive what is relevant and, at the same time, reduce the noise in the sensory streams using gaze stabilization, for example.

### v2

@@ IV-C Real robot – bimanual task
IV-D Real robot – Interactive game demonstration
To demonstrate the operation of HARMONIOUS at scale and in a genuine unstructured human-robot interaction scenario, we taught the robot to play a children’s board game called Bubbles [^1] A description of the game is available at https://www.piatnik.com/en/games/board-games/children-games/bubbles. For our purposes, we simplified the game, with only six cards, three dice, only the human rolling the dice, and without grasping the cards (only pointing at the selected one). The robot recognizes the cards but is told the numbers on the dice after the roll (as these are too small to be recognized).. Based on the numbers that appear on the dice and their colors, the players—the robot and the human in our case—compete to pick the card on the table where the size of the colored bubbles corresponds to the colors of the dice, ordered by the numbers – see Fig. 1 and video S5 in Multimedia Materials. This game was chosen because there is a shared physical space with unstructured interaction. The robot processes the visual inputs and recognizes the individual cards on the table. After the roll of the dice, the player who first places his hand over the right card wins the round. After choosing the appropriate card, the robot selects which hand to use and initiates the reach movement. At all times and in real time, visual, proximity, and tactile streams are simultaneously processed, spawning obstacles if appropriate and dynamically adding constraints to the robot controller in order to warrant the safety of the human player at all times. The game itself is rather challenging for the human player and hence in order to compete with the robot, he often cheats by pushing the robot physically away or by blocking certain volumes of the workspace by placing his hands over it. This was a deliberate choice to put HARMONIOUS at test.
The controller runs in dual-arm mode during the game. Depending on the position of the card in every round, the left or right arm is chosen and set as “primary’ (and hence its position reaching target is set as a constraint). Additional constraints were added to the optimization problem by adding static obstacle points (see Sec. III-I) to prevent the robot from colliding with the table. We prepared a high-level Python script to play the game and used the gaze controller [60] to always look at the table. The program loop starts with the processing of rolled dice (numbers sent to the program in the command line) and the current RGB-D image, followed by the detection of cards on the table, the decision of the correct card and computing its 3D position. Then one arm approaches the card position, while the other arm goes to a specific position outside the field of view of the RGB-D camera. After the reach approach (successful or not), the approaching arm returns to a specific position to clear the view for the next round.
@@ IV-D Real robot – Interactive game demonstration
V Conclusion, Discussion and Future work
@@
We have developed a real-time reactive motion control system for upper body control of a dual-arm robot with a common torso. The proposed solution (HARMONIOUS) is designed to enable safe and effective interaction with humans in close proximity. In the absence of obstacles, we systematically tested our controller’s reach ability on a grid of position and orientation targets in the workspace and compared it to NEO [2] and React [3, 4], demonstrating superior performance. We found that assigning different weights to position and orientation tasks is advantageous (unlike in [2]). Preferring arm over torso movements by adjusting weights in the minimization leads to more natural-looking movements. Uniquely to our solution, we commanded both arms and the torso of the iCub humanoid robot (in total $17$ DoFs) and demonstrated both the setup where every arm has a separate task and one, set as “primary arm”, commands the torso joints, and a bimanual task where the hands hold an object between them. Kinematic singularities are handled through velocity damping and natural kinematic postures are rewarded in the optimization problem formulation. Smooth human-like motions with a bell-shaped Cartesian velocity profile are generated using biologically inspired local trajectory sampling. Overall, several components act synergistically to produce human-like naturally looking movements.
The solution shows the online dynamic incorporation of obstacles perceived through three different sensory modalities and their on-the-run embedding as whole-body motion constraints into a motion controller with human-like characteristics. Two of these modalities, vision and proximity, sense at a distance, “pre-collision”, and the other one—tactile sensors on large areas of the robot body ($2000$ individual sensors)—is providing contact or post-collision information. The performance of HARMONIOUS at scale was finally demonstrated in unstructured physical human-robot interaction while playing a game.
The merit of HARMONIOUS lies in the ability to perceive objects in the environment, relate them to the complete surface of the robot body, and, when appropriate, transform them into constraints for the motion controller. Information about obstacles from all three modalities—sometimes overlapping, at other times complementary—was processed and transformed into a unified representation dynamically generating constraints for the whole-body motion controller.
The number or type of sensors used in this work is by no means a requirement for the described system to work. The solution will work with less dense coverage of the perirobot space, covered by only one sensor type, for example. However, dense and multimodal coverage is advantageous for the safety of the interaction. At the same time, additional sensor types could be easily incorporated. Visual obstacles and the corresponding constraints will work for any range-based sensor. Physical contact is implemented here through tactile sensors and acts as a constraint along the contact normal. If a contact force vector was available, it could be incorporated preserving the information about the direction of the constraint (as is currently done for the visual obstacles). One limitation of our current implementation is that the obstacles perceived by one sensory modality are combined into a single motion constraint per robot body part. This was an implementation decision on the obstacle processing layer. Alternatively, multiple constraints from a single modality could be preserved, transformed into constraints, and fed into the motion controller. Or, the information about obstacles from mutliple modalities could be first combined, preserving more information about the context (see contact hypotheses fusion [49]) and only then transformed into motion constraints.
HARMONIOUS is a versatile and extendable whole-body motion controller which, however, remains at the kinematic level. The problem could be approached using robot dynamics and formulated at the level of forces and torques. This may be advantageous (solving inverse dynamics rather than inverse kinematics) and ultimately more general, but at the same time it would require accurate identification of the robot inertial parameters. In general, a “dynamics controller” constitutes an alternative, different approach. With HARMONIOUS we have shown the potential and versatility of the kinematic approach, which in our view is simpler to deploy in any robot where a kinematic model and position or velocity control are available.
The problem formulation is highly modular and both the minimization criteria (e.g. motivating preferred postures) or constraints to the quadratic program (e.g. obstacles) can be easily removed or different ones added. For example, additional task constraints were added to the problem formulation for the interactive game scenario. On the other hand, our solution guarantees strict task prioritization only between constraints (e.g., obstacle avoidance) and objective function minimization (e.g., target reaching) and not between individual tasks. To enforce strict task priority, a framework of prioritized QP-based kinematic control with a null space projection can be incorporated [61].
@@
Our immediate future work involves adding active gaze control such that the robot can better perceive what is relevant and, at the same time, reduce the noise in the sensory streams using gaze stabilization, for example.